### v2

# Wenxuan Zhou Google DeepMind Shujian Zhang Google DeepMind Brice Magdalou CEE-M, Montpellier U. John Lambert Google DeepMind Ehsan Amid Google DeepMind Richard Nock Google Research Andrew Hard Google DeepMind

Abstract
In this paper, we show that direct preference optimization (DPO) is a very specific form of a connection between two major theories in the ML context of learning from preferences: loss functions (Savage) and stochastic choice (Doignon-Falmagne and Machina). The connection is established for all of Savage’s losses and at this level of generality, (i) it includes support for abstention on the choice theory side, (ii) it includes support for non-convex objectives on the ML side, and (iii) it allows to frame for free some notable extensions of the DPO setting, including margins and corrections for length. Getting to understand how DPO operates from a general principled perspective is crucial because of the huge and diverse application landscape of models, because of the current momentum around DPO, but also – and importantly – because many state of the art variations on DPO definitely occupy a small region of the map that we cover. It also helps to understand the pitfalls of departing from this map, and figure out workarounds.
keywords:
Direct Preference Optimization, Proper Loss Functions, Stochastic Choice Theory
1 Introduction
Direct Preference Optimization (DPO) introduced a trick to bypass reward modelling in RLHF, which trains reward function $r_{\bm{\omega}}$ then policy $\pi_{\bm{\theta}}$ with parameters $\bm{\theta}$, by introducing an explicit direct link $r_{\bm{\theta}}=f(\pi_{\bm{\theta}})$ that allows to skip learning the reward function parameters ($\bm{\omega}$) (Rafailov et al., 2023). While the technique has met with experimental success, a trove of papers have elaborated on the DPO framework to generalize its functional parts (Alfano et al., 2025; Azar et al., 2024; Chen et al., 2025; Cui et al., 2024; Gu et al., 2024; Gupta et al., 2025; Hong et al., 2024; Lee et al., 2025; Li et al., 2025; Lu et al., 2024; Meng et al., 2024; Park et al., 2024; Ramesh et al., 2024; Yuan et al., 2023; Zhao et al., 2025) (and many others). This is understandable: there is lots of apparent freedom from this standpoint, with two directions to explore, loss functions — two in the DPO pipeline — and choice theory. It is an understatement to say that there is a push to explore those dimensions (Zhang et al., 2025), but the risks faced without a principled understanding of how the connection works can be consequential: the core element of any learning algorithm, the objective to be optimized, involves design elements from both choice theory and loss functions. No work has carried out this principled grounding of the connection. This is our focus.
We show that DPO is a very particular case of intertwining two foundational theories for loss functions and stochastic choice in the ML context of learning from preferences. The former is Savage’s properness (Savage, 1971). The latter involves elements from Doignon-Falmagne’s stochastic choice theory (Doignon and Falmagne, 1974) as simplified by Krantz, Luce, Suppes and Tversky in their book (Krantz et al., 1989), and from Machina on lotteries (Machina, 1985; Machina and Siniscalchi, 2014).
These many design elements lead to a generalization making the most of the connection since we encompass all of properness on Savage’s side (regardless of optional properties like symmetry, differentiability, etc). We also encompass all of the modelling’s power on Krantz, Luce, Suppes and Tversky’s side. At this level of generality and through the connection, we are able to cover on the choice theory side the case where users can abstain, provided abstention has a particular structure among alternatives; and we are able to cover non-convex losses in the final ML-optimized objective, provided they are strictly monotonic. If abstaining is forbidden, which is the setting of DPO and most of its current ML generalizations, we can reduce loss choices to symmetric losses on Savage’s side (an example of which is the log-loss, strictly proper, and its dual convex conjugate, the logistic loss, strictly convex) and Debreu models for the choice theory side (Debreu, 1958) (an example of which is Bradley-Terry-Luce’s). Notably, our level of generalization is able to support "for free” important extensions of DPO that have been so far essentially ad-hoc: margins / home advantages (Meng et al., 2024) and corrections for length (Meng et al., 2024; Gu et al., 2024; Gupta et al., 2025; Park et al., 2024; Chen et al., 2025).
This is an important task because DPO was designed with the objective to simplify RLHF and getting "above" DPO is mandatory to improve results by getting more freedom on reward shapes, trajectories and preference behaviours (Gupta et al., 2025), all of which needs to be done while keeping DPO’s design tricks supporting simplification. What is striking, at the level of the panorama that we reach, is that it groups many DPO variants in a pretty small part, while the rest of it remains largely terra incognita. We provide several ways to navigate broad parts of this panorama: the most general which involves a loss eventually non-convex, and a particular one, convenient when all or part of the functional framework needs to be learned (Alfano et al., 2025), which trades expensive properness checks for a cheap convexity knob.
Our work is also interesting for what may depart from this map, to understand the pitfalls and figure out workarounds. One perhaps unexpected pitfall comes from the RLHF/DPO inherited "gold standard" for the relationship $r(x,y)=f(\pi({\color[rgb]{1,0,0}y}|x))$, whereby $f$ depends only on the action / prompt $y$ involved in the reward. We show that there is only one proper loss for which such a representation is possible, DPO’s KL divergence. Our proof eliminates assumptions sprinkled on seven decades of related results (McCarthy, 1956; Savage, 1971; Bernardo, 1979; Dawid, 2007): differentiability, the identity of pointwise losses, restrictions on the target, etc. . Without care, changing the KL divergence while keeping properness can bring $r(x,y)=f(\mbox{{many} }\pi({\color[rgb]{1,0,0}y^{\prime}}|x))$, potentially ladening the objective with a computational burden contrasting with DPO’s simplicity. Fortunately, we show that there exists a very simple workaround in the "gold standard", keeping properness with much freedom in loss design.
To preserve readability, all proofs are given in an appendix.
2 Definitions
We adopt many definitions from Rafailov et al. (2023). States / contexts / prompts are denoted by $x_{(.)}\in\mathcal{X}$, actions / alternatives / answers by $y_{(.)}\in\mathcal{Y}$. Let $m$ be the size of $\mathcal{X}$ and $n$ the size of $\mathcal{Y}$. We sometimes enumerate the elements of $\mathcal{Y}$ in $\mathbb{N}$ as $y_{1},y_{2},...,y_{n}$. $\Delta_{\mathcal{Y}}$ being the set of distributions over $\mathcal{Y}$, a policy $\pi$ is an element of $\Delta_{\mathcal{Y}}^{\mathcal{X}}$, associating to each state a distribution over actions. There exists a reference (pretrained) policy, $\pi_{\mbox{\tiny{ref}}}$. We denote $n\stackrel{{\scriptstyle\mathrm{.}}}{{=}}|\mathcal{Y}|$ and $m\stackrel{{\scriptstyle\mathrm{.}}}{{=}}|\mathcal{X}|$ and abbreviate the simplex $\Delta_{n}$ for short.
2.1 Properness
A loss function is a function $L(\bm{p},\bm{q}):\Delta_{n}\times\Delta_{n}\rightarrow\mathbb{R}$.
Definition 1.
Fix $\bm{p}\in\Delta_{n}$. Loss $L$ is $\bm{p}$-proper iff $L(\bm{p},\bm{q})\geq L(\bm{p},\bm{p}),\forall\bm{q}\in\Delta_{n}$. It is proper iff it is $\bm{p}$-proper for any $\bm{p}\in\Delta_{n}$.
Strict properness happens if the inequality is an equality only for $\bm{q}=\bm{p}$. Properness encourages guessing the ground truth. Savage proposed to use an appealing form for $L$:
$L(\bm{p},\bm{q})$ $\stackrel{{\scriptstyle\mathrm{.}}}{{=}}$ $\bm{p}^{\top}\bm{\ell}(\bm{q})=\mathbb{E}_{i\sim\textsc{cat}(\bm{p})}[\ell_{i}(\bm{q})],$ (1)
where "cat" denotes a categorical distribution with parameter $\bm{p}$. Properness is intensional (Walder and Nock, 2020) and can be tricky to operate if (parts of) the loss itself need to be learned (Alfano et al., 2025; Gupta et al., 2025). Fortunately, there is a characterization of proper losses that makes things a lot easier. A Bregman divergence $D_{\phi,\bm{H}}(\bm{u}\|\bm{v})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\phi(\bm{u})-\phi(\bm{u})-(\bm{u}-\bm{v})^{\top}\bm{H}(\bm{v})$ is defined for convex $\phi$ and a selection of its subdifferential, $\bm{H}\in\partial\phi$. Bregman divergences have a second form, introduced in information geometry by Amari and Nagaoka (2000) and later introduced in ML under the name Fenchel-Young losses (Blondel et al., 2020), whereby $D_{\phi}(\bm{u}\|\bm{v}^{*})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\phi(\bm{u})+\phi^{\star}(\bm{v}^{*})-\bm{u}^{\top}\bm{v}^{*}$, where $\varphi^{\star}(\bm{v})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\sup_{\bm{t}}\bm{t}^{\top}\bm{v}-\varphi(\bm{t})$ is the convex conjugate. This Amari-Bregman-Fenchel (ABF) representation keeps implicit the dependence on the subgradient, but it is right there since $\bm{v}^{*}=\bm{H}(\bm{v})$ for some $\bm{H}\in\partial\phi$. All proper losses can be represented as Bregman divergences, and reciprocally every Bregman divergence "encodes" a proper loss (Williamson et al., 2016).
2.2 Stochastic choice with abstention
We build our setting upon the simplified treatment of the seminal work of Doignon and Falmagne (1974) by Krantz et al. (1989, Chapter 17), to which we inject Machina’s well known lottery approach to objective uncertainty (Machina, 1985; Machina and Siniscalchi, 2014). For that reason, we shall use the acronym KLST to name our framework.
Expandability defines the way we include simple lotteries, that is, binary choice lotteries. The lottery choice $({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}$ ($\alpha\in[0,1]$) is revealed through randomization, which displays ${\color[rgb]{1,0,0}y_{1}}\in\mathcal{Y}$ with probability $\alpha$ and ${\color[rgb]{1,.5,0}y_{2}}\in\mathcal{Y}$ with probability $1-\alpha$. Stochastic choices involving lotteries are naturally computed as
$\left\{\begin{array}[]{ccc}p(({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}\succ L|x)&=&\alpha\cdot p({\color[rgb]{1,0,0}y_{1}}\succ L|x)+(1-\alpha)\cdot p({\color[rgb]{1,.5,0}y_{2}}\succ L|x)\\ p(L\succ({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}|x)&=&\alpha\cdot p(L\succ{\color[rgb]{1,0,0}y_{1}}|x)+(1-\alpha)\cdot p(L\succ{\color[rgb]{1,.5,0}y_{2}}|x)\end{array}\right.$. (4)
Putting a density on $\alpha$ yields a sampling process totally compatible with the triplet sampling process $(x,y,y^{\prime})$ against which preference optimization is usually done. If deemed preferable or necessary (if e.g. picking $\alpha$ does not follow a stochastic process), one could also consider having the user face lotteries directly in the choices to make. Our framework can still work as is provided one makes the additional "Reduction of Compound Lotteries Axiom" (Machina and Siniscalchi, 2014). Fix $\alpha\in[0,1]$. Let $\mathcal{Y}^{\alpha}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\{(yy^{\prime})_{\alpha}:y,y^{\prime}\in\mathcal{Y}\}$ be the set of available lotteries. An important point to remember is that lottery $(yy)_{\alpha}$ is equivalent to alternative $y$ because any choice probability involving $(yy)_{\alpha}$ breaks down to the one involving only $y$ via (4). Hence $\mathcal{Y}^{\alpha}$ "contains" $\mathcal{Y}$. We say that choice probabilities are expandable to denote that the support of all actions $\mathcal{Y}$ is extended to cover all lotteries.
Local choice structure Fix $\alpha\in[0,1]$. Each $x\in\mathcal{X}$ defines an undirected graphs $G^{\alpha}_{x}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\mathcal{Y}^{\alpha},\mathcal{E}^{\alpha}_{x})$ where edges are as follows:
$(L,L^{\prime})\in\mathcal{E}^{\alpha}_{x}$ $\Leftrightarrow$ $p(L\succ L^{\prime}|x)+p(L^{\prime}\succ L|x)=1.$ (5)
Edges encode zero abstention (ZA) in a choice between lotteries. We also introduce the shorthand
$L\triangleright_{x}L^{\prime}$ $\Leftrightarrow$ $p(L\succ L^{\prime}|x)\geq 1/2$ (6)
($\alpha$ implicit), indicating that lottery $L$ is preferred (P) over $L^{\prime}$ given $x$. In an undirected graph, a wedge $\{a,b,c\}$ is a sequence of three vertices that form a chain as $(a,b),(b,c)$. Our definition of a local choice structure, which extends to lotteries that of Krantz et al. (1989, Chapter 17), follows.
Definition 2.
Lotteries yield a Local Choice Structure (LCS) iff the following three properties are satisfied for any $\alpha\in(0,1)$:
  • Bearability: $(L,L)\in\mathcal{E}^{\alpha}_{x},\forall L\in\mathcal{Y}^{\alpha}$;
[…]
Monotonicity To wrap up our model, we need a property among relative preferences (Krantz et al., 1989).
Definition 3.
Lotteries satisfy monotonicity iff there exists $\alpha\in(0,1)$ such that the following holds. For any ${\color[rgb]{1,0,0}L_{1}},{\color[rgb]{1,.5,0}L_{2}},{\color[rgb]{1.0,0.88,0.21}L_{3}},{\color[rgb]{0,0.8046875,0}L_{4}},{\color[rgb]{0,1,1}L_{5}},{\color[rgb]{.5,0,.5}L_{6}}$ in $\mathcal{Y}^{\alpha}$, if (i) $\{{\color[rgb]{1,0,0}L_{1}},{\color[rgb]{1,.5,0}L_{2}},{\color[rgb]{1.0,0.88,0.21}L_{3}}\}$ is a triangle in $\mathcal{E}_{x}^{\alpha}$, (ii) $\{{\color[rgb]{0,0.8046875,0}L_{4}},{\color[rgb]{0,1,1}L_{5}},{\color[rgb]{.5,0,.5}L_{6}}\}$ is a wedge in $\mathcal{E}_{x}^{\alpha}$, (iii) $p({\color[rgb]{1,0,0}L_{1}}\succ{\color[rgb]{1,.5,0}L_{2}}|x)\geq p({\color[rgb]{0,0.8046875,0}L_{4}}\succ{\color[rgb]{0,1,1}L_{5}}|x)$ and (iv) $p({\color[rgb]{1,.5,0}L_{2}}\succ{\color[rgb]{1.0,0.88,0.21}L_{3}}|x)\geq p({\color[rgb]{0,1,1}L_{5}}\succ{\color[rgb]{.5,0,.5}L_{6}}|x)$, then
$(({\color[rgb]{0,0.8046875,0}L_{4}},{\color[rgb]{.5,0,.5}L_{6}})\in\mathcal{E}_{x}^{\alpha})\wedge(p({\color[rgb]{1,0,0}L_{1}}\succ{\color[rgb]{1.0,0.88,0.21}L_{3}}|x)\geq p({\color[rgb]{0,0.8046875,0}L_{4}}\succ{\color[rgb]{.5,0,.5}L_{6}}|x)).$ (7)
The large number of parameters may create an impression of complexity which hides a simple core message: condition (iii) may be read $\color[rgb]{1,0,0}L_{1}$} is preferred to $\color[rgb]{1,.5,0}L_{2}$} more than $\color[rgb]{0,0.8046875,0}L_{4}$} is preferred to $\color[rgb]{0,1,1}L_{5}$}" and condition (iv) is $\color[rgb]{1,.5,0}L_{2}$} is preferred to $\color[rgb]{1.0,0.88,0.21}L_{3}$} more than $\color[rgb]{0,1,1}L_{5}$} is preferred to $\color[rgb]{.5,0,.5}L_{6}$}". The right consequence in (7) is then intuitive: $\color[rgb]{1,0,0}L_{1}$} is preferred to $\color[rgb]{1.0,0.88,0.21}L_{3}$} more than $\color[rgb]{0,0.8046875,0}L_{4}$} is preferred to $\color[rgb]{.5,0,.5}L_{6}$}". The left consequence in (7) imposes that the wedge on condition (ii) be in fact a triangle in $\mathcal{E}_{x}^{\alpha}$, imposing no abstention when choosing $\color[rgb]{0,0.8046875,0}L_{4}$} vs $\color[rgb]{.5,0,.5}L_{6}$} (which is natural considering (i) and (ii)). Importantly, monotonicity needs to be defined only for one non-trivial $\alpha$ ($\in(0,1)$), which is different from the requirements of a LCS.
KLST structure
We are now ready to state  the definition of a KLST structure.
Definition 4.
The choice probabilities $p(y\succ y^{\prime}|x)$ have a KLST structure iff they are expandable and the resulting lotteries (i) yield a local choice structure and (ii) satisfy monotonicity, $\forall x\in\mathcal{X}$.
If we were to impose $G_{x}^{\alpha}$ to be a clique, i.e. to never abstain to pick a choice, then this setting could be substantially simplified to the seminal setting of (Debreu, 1958), encompassing as particular case the well-known Bradley-Terry-Luce model. Additional details about KLST are in the Appendix, Section VI.
3 Preference optimization with KLST and Savage
Figure 1: DPO’s recipe abstracted, where we indicate the theoretical object allowing to pick the key component in $f,h,I$: either loss function or choice theory. Parameter $\color[rgb]{0,0.8046875,0}\bm{\theta}$} are those of the model to be learned. See text for details.
3.1 Direct Preference Optimization (DPO)
We first briefly review preference optimization as it is carried out in the paper (Rafailov et al., 2023). Figure 1 summarizes the key steps of DPO alongside the theoretical objects (loss and stochastic choice theory) allowing to cherry pick DPO’s components. We summarize these components, putting in blue their remarkable specific parts.
Step (1) establishes the link which later avoids reward fitting. Function $f$ is computed from the solution of the following problem [^1] This is equivalent to computing the full policy over the expectation over $x$.:
$\pi(.|x)$ $=$ $\arg\max_{\pi^{\prime}(.|x)\in\Delta_{n}}\mathbb{E}_{y\sim\pi^{\prime}(.|x)}[r(x,y)]-D_{{\color[rgb]{0,0,1}\textsc{kl}}}(\pi^{\prime}(.|x)\|\pi_{\mbox{\tiny{ref}}}(.|x)),\forall x\in\mathcal{X},$ (8)
Where kl denotes the KL divergence, a Bregman divergence. Crucially for the computational benefits of the approach, the Lagrange multiplier which hides the cumulant function disappears from the difference of two rewards.
Step (2) links the probability of choosing $y$ against $y^{\prime}$ given $x$ to their corresponding rewards:
$p_{\color[rgb]{0,0.8046875,0}\theta}(y\succ y^{\prime}|x)$ $\stackrel{{\scriptstyle\mathrm{.}}}{{=}}$ ${\color[rgb]{0,0,1}\frac{1}{1+\exp(-({\color[rgb]{0,0,0}r_{\color[rgb]{0,0.8046875,0}\theta}(y,x)-r_{\color[rgb]{0,0.8046875,0}\theta}(y^{\prime},x)}))}},$ (9)
so $h$ in Figure 1 is the sigmoid function.
Step (3) The final loss to be optimized in DPO is then a function of the policy to be learned,
$I(p_{\color[rgb]{0,0.8046875,0}\theta})$ $=$ $\mathbb{E}_{(x,y,y^{\prime})\sim\mathcal{D}}\left[{\color[rgb]{0,0,1}\psi}\left({\color[rgb]{0,0,1}\log}\frac{\pi_{\color[rgb]{0,0.8046875,0}\theta}(y|x)}{\pi_{\mbox{\tiny{ref}}}(y|x)}-{\color[rgb]{0,0,1}\log}\frac{\pi_{\color[rgb]{0,0.8046875,0}\theta}(y^{\prime}|x)}{\pi_{\mbox{\tiny{ref}}}(y^{\prime}|x)}\right)\right],$ (10)
where ${\color[rgb]{0,0,1}\psi(}z{\color[rgb]{0,0,1})}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}{\color[rgb]{0,0,1}\log(1+\exp(-}z{\color[rgb]{0,0,1}))}$ is the logistic loss. It is remarkable that all the blue parts are tightly linked: the KL divergence is the Bregman divergence of the log-loss, the sigmoid is its inverse canonical link and the logistic loss is its dual convex surrogate (Nock and Nielsen, 2008; Reid and Williamson, 2010). Such a remarkable connection is not fortuitous: the rest of this section generalizes it.
3.2 Generalization of step (1)
The central trick of DPO relies on creating the link $r_{\color[rgb]{0,0.8046875,0}\theta}=f(\pi_{\color[rgb]{0,0.8046875,0}\theta})$ (Figure 1) by solving a generalized form of (8) as $\pi(.|x)=\arg\max_{\pi^{\prime}(.|x)}J_{\textsc{kl},r}(\pi^{\prime}(.|x))$. Let us abstract $J$ as:
$J_{R,r}(\pi(.|x))\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathbb{E}_{y\sim\pi(.|x)}[r(x,y)]-R(\pi(.|x)\|\pi_{\mbox{\tiny{ref}}}(.|x)),$ (11)
where $R$ is the regret of a function $Q(\pi(.|x)\|\pi_{\mbox{\tiny{ref}}}(.|x))$ that we aim to elicit. The regret being the difference between $Q$ and its infimal value (assumed finite wlog), it is non-negative. Such a general formulation is ubiquitous in ML. The question we address is: what form should have $Q$ ? There is a simple condition that imposes a simple analytical structure generalizing the KL divergence. Suppose we have learned
$\pi_{*}(.|x)$ $\stackrel{{\scriptstyle\mathrm{.}}}{{=}}$ $\arg\max J_{R,r}(\pi(.|x)),$
and we want to carry out a second stage of training. We must ensure that $\pi_{*}(.|x)$ is the best initialization, just like $\pi_{\mbox{\tiny{ref}}}(.|x)$ was for the first stage, regardless of the rewards to be involved. This leads to a specific request on the objective:
$\pi(.|x)$ $\in$ $\arg\min_{\pi^{\prime}(.|x)}Q(\pi(.|x)\|\pi^{\prime}(.|x)).$ (12)
In the jargon of Savage, this means imposing $Q$ be proper.
Theorem 5.
Suppose $Q$ proper. Then, using notations from (12),
$Q(\pi(.|x)\|\pi^{\prime}(.|x))$ $=$ $K(\pi(.|x))+D_{\phi,G}(\pi(.|x)\|\pi^{\prime}(.|x)),$ (13)
where $D_{\phi,G}$ is a Bregman divergence. More specifically, we can show $\phi(\bm{p})=-L(\bm{p},\bm{p})$ for some proper loss $\ell$ (1) and can pick $G_{i}=-\ell_{i}$.
Proof in Appendix, Section VII.1. Note that $G_{i}$ is allowed to depend not just on $y_{i}$. It comes from Williamson et al. (2016, Proposition 7) that $Q(\pi(.|x)\|\pi(.|x))=K(\pi(.|x))$, so we get the expression of the regret in (11),
$R(\pi(.|x)\|\pi^{\prime}(.|x))$ $\stackrel{{\scriptstyle\mathrm{.}}}{{=}}$ $Q(\pi(.|x)\|\pi^{\prime}(.|x))-Q(\pi(.|x)\|\pi(.|x))$
$=$ $D_{\phi,G}(\pi(.|x)\|\pi^{\prime}(.|x)).$
3.3 Generalization of step (2)
[…]
Theorem 6.
Suppose $p(y\succ y^{\prime}|x)$ has a KLST  structure. Then there exists
  • a strictly increasing function $F$ with $\mathrm{Im}F\subseteq[0,1]$ and such that $F(-z)+F(z)\leq 1,\forall z\in\mathrm{dom}F$;
  • a function $u(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$,
such that
$p(y\succ y^{\prime}|x)$ $=$ $F(u(x,y)-u(x,y^{\prime})).$ (14)
Proof in Appendix, Section VII.2. One can show that we get from KKT conditions in the optimization of (11) with $Q$ as in (13) the relationship (indexed notations follow Theorem 5) for any $i\in[n]$,
$r(x,y_{i})+0_{\pi(y_{i}|x)>0}\cdot\mu_{y_{i}}$ $=$ $\lambda+G_{i}(\pi(.|x))-G_{i}(\pi_{\mbox{\tiny{ref}}}(.|x)),\quad\bm{G}\in\partial\phi.$ (15)
Here, $\lambda$ is the Lagrange multiplier enforcing normalization of $\pi(.|x)$, $\mu_{y_{i}}\geq 0$ is the one for the constraint of non-negativity for $\pi(y_{i}|x)$; finally, for a predicate $\Pi$, $0_{\Pi}$ takes value 0 if $\Pi$ is true and 1 otherwise. If we plug $u(x,y_{i})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}r(x,y_{i})+0_{\pi(y_{i}|x)>0}\cdot\mu_{y_{i}}$ (if all policy values are >0>0 such as in DPO, this is just the reward function) in (15), then we remark for any $x,y,y^{\prime}$,
$u(x,y_{i})-u(x,y_{j})$ $=$ $G_{i}(\pi(.|x))-G_{i}(\pi_{\mbox{\tiny{ref}}}(.|x))-G_{j}(\pi(.|x))+G_{j}(\pi_{\mbox{\tiny{ref}}}(.|x)),$ (16)
so just like in DPO, the expression of $p(y_{i}\succ y_{j}|x)$ in Theorem 6 does not depend on the (computationally expensive) $\lambda$. There remains to see how the final loss $I$, in Step (3), relates to Savage’s theory in our broad setting.
3.4 Generalization of step (3) with a canonical connection
Having to minimize a function $I$ that depends on probabilities, Savage’s properness framework is again the instrument of choice. The state of the art focuses on the minimization of $I(\pi_{\bm{\theta}})=\mathbb{E}_{(x,y,y^{\prime})}\left[\ell(p_{\bm{\theta}}(y_{i}\succ y_{j}|x))\right]$ for some $\ell:[0,1]\rightarrow\mathbb{R}$. After Steps (1) and (2), this objective becomes
$I(\pi_{\bm{\theta}})$ $=$ $\mathbb{E}_{(x,y_{i},y_{j})}\left[\ell\circ F\left(G_{i}(\pi_{\bm{\theta}}(.|x))-G_{i}(\pi_{\mbox{\tiny{ref}}}(.|x))-G_{j}(\pi_{\bm{\theta}}(.|x))+G_{j}(\pi_{\mbox{\tiny{ref}}}(.|x))\right)\right],$ (17)
for policy parameters $\bm{\theta}$ and all other parameters defined in Theorems 5 and 6. There are two important things to note here:
  • this problem is equivalent to a binary classification problem where $(x,y_{i},y_{j})$ is positive if say $y_{i}\succ y_{j}$ given xx" and negative if $y_{j}\succ y_{i}$ given xx" (Sun et al., 2025) so the proper loss is scalar valued, which we note $L(p,p)=p\ell_{1}(p)+(1-p)\ell_{0}(p)$ after Reid and Williamson (2010). The indexing in $\{0,1\}$ is a convention departing from our indexing in 1,2,,n1,2,...,n (Definition 1): "0" usually refers to the so-called "negative class" and "1" to the "positive class". Such a loss is noted $(\ell_{0},\ell_{1})$;
  • most DPO generalizations have so far considered a restriction to symmetric losses (Nock and Nielsen, 2008), for which we have $\ell_{0}(p)=\ell_{1}(1-p)$ and (17) implicitly assumes the usual convention $\ell\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{1}$.
We now tackle the problem from the most general standpoint and drop the symmetry assumption. What are conditions for which (17) bears relevant properties in the context of Savage’s theory, given that the arguments of $\ell\circ F$" in (17) are reals and not probabilities ?
Theorem 7.
For any strictly proper loss $(\ell_{0},\ell_{1})$, denote $\phi(p)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}-L(p,p)$. Letting $H(p)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{0}(p)-\ell_{1}(p)$, we have (i) $H\in\partial\phi$ and (ii) $\ell_{0}\circ H^{-1}(x)=\phi^{\star}(x)$. If furthermore the loss is symmetric, then we also have (iii) $\ell\circ H^{-1}(x)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{1}\circ H^{-1}(x)=\phi^{\star}(-x)$.
Proof in Appendix, Section VII.3. We now easily unveil the general form for (17), which is obtained either using the convention $\ell_{1}\circ H^{-1}(x)$ (only for symmetric losses) or the more general form $\ell_{0}\circ H^{-1}(-x)$ for general losses, and thus bonding loss and social choice theories.
Corollary 8.
For any strictly proper loss $(\ell_{0},\ell_{1})$, suppose we pick $\color[rgb]{1,0,0}F=H^{-1}$} in Theorem 6 where $H$ is as in Theorem 7 and let $\ell\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{0}$ in (17). Then (17) simplifies to:
$I(\pi_{\bm{\theta}})$ $=$ $\mathbb{E}_{(x,y_{i},y_{j})}\left[\psi\left(-[G_{i}(\pi_{\bm{\theta}}(.|x))-G_{i}(\pi_{\mbox{\tiny{ref}}}(.|x))-G_{j}(\pi_{\bm{\theta}}(.|x))+G_{j}(\pi_{\mbox{\tiny{ref}}}(.|x))]\right)\right].$ (18)
with $\psi\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\phi^{\star}$.
The red part in Corollary 8 establishes the "connection" between choice theory and loss functions: $F$ gives choice probabilities in KLST, $H$ maps to the loss via the subdifferential in Savage. We call it a canonical connection because the choice function $F$ is directly related to a canonical link of the loss, $H$.
Remark 9.
If $F=H^{-1}$, condition $F(-z)+F(z)\leq 1$ (Theorem 6) is equivalent to
$\ell_{0}\left(\frac{1}{2}+\delta\right)+\ell_{0}\left(\frac{1}{2}-\delta\right)$ $\geq$ $\ell_{1}\left(\frac{1}{2}+\delta\right)+\ell_{1}\left(\frac{1}{2}-\delta\right),\forall\delta\in(0,1/2).$
While it is met for symmetric losses, if an asymmetric loss $(\ell_{0},\ell_{1})$ does not meet those requirements, then the "flipped" loss $(\ell_{1}(1-p),\ell_{0}(1-p))$ does.
After DPO, preference optimization seems to have essentially focused on symmetric losses. Even in the most general case, such losses on real-valued predictions have a long story in binary classification. The link between (18) and Savage’s properness is indeed simple: it is equivalent from the optimization standpoint to a proper loss, where the connection between real values and probabilities is done via the canonical link of the loss (Nock and Nielsen, 2009, Lemma 2). Hence, the minimization of (18) occurs in disguise in the properness framework of Savage, which of course was the objective. While being convex, $\phi^{\star}$ has also additional interesting properties, such as being 1-Lipschitz and strictly monotonic.
Remark 10.
(the modelling "price" for never abstaining) Forbidding abstention yields to stochastic choice models pioneered by Debreu, enforcing F(z)+F(z)=1F(z)+F(-z)=1. Here comes an interesting fact: the simplification using the convex conjugate in Corollary 8 and commonly used in the state of the art DPO-style analysis happens only if the strictly proper loss chosen in Theorem 7 is symmetric.
3.5 Further generalization of step (3) with a composite connection
In the canonical connection used in Corollary 8, a strictly proper loss $(\ell_{0},\ell_{1})$ defines both $\ell$ and $F$ in the objective $I(\pi_{\bm{\theta}})$ (17), $\ell$ being $\ell_{0}$ and $F$ the inverse of a canonical link of the loss. This yields a broad family of losses that rely on the convex surrogate of the loss, $\phi^{\star}(-z)$, but as we said, they all share the same blueprint (1-Lipschitz, etc.). Let us strive for the maximum generality in covering "acceptable" losses for the objective $I(\pi_{\bm{\theta}})$ (17): what is the set of functions $\psi:\mathbb{R}\rightarrow\mathbb{R}$ that could be used in lieu of $\phi^{\star}$ while still guaranteeing the existence of (i) a choice function $F$ meeting the conditions of Theorem 6, (ii) a strictly proper loss $(\ell_{0},\ell_{1})$, and such that $\psi=\ell_{0}\circ F$ ? (we refer to such a connection as being "composite"). The result, whose proof is constructive, shows a very broad result. First, Lipschitzness is (predictably) optional, but convexity is also optional. Second, for any applicable $\psi$, any $F$ meeting the conditions of Theorem 6 can work.
Theorem 11.
For any strictly increasing function $\psi:\mathbb{R}\rightarrow\mathbb{R}$, any strictly increasing function $F:\mathbb{R}\rightarrow[0,1]$, there exists a strictly proper loss $(\ell_{0},\ell_{1})$ such that
$\psi(z)$ $=$ $\ell_{0}\circ F(z).$ (19)
Proof in Appendix, Section VII.4. The trick that expedites the proof comes from a result later exposed in Section 4.3: (20) builds a proper loss from a convex function, but it can be read as well as an ODE giving the convex function from a function (a partial loss), from which we build the complete loss and check it is indeed strictly proper.
Note that we cannot have a result in which we would instead pick any strictly proper loss and then get $F$ meeting the conditions of Theorem 6. Here is a simple counterexample: pick $\psi(z)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\exp(z)$ (gives the exponential loss), and $\ell_{0}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}-\log(1-p)$ (the log-loss). We get $F=\ell_{0}^{-1}\circ\psi=1-\exp(-\exp(z))$, which violates the condition $F(-z)+F(z)\leq 1,\forall z\in\mathbb{R}$ (F(0)=1(1/e)>1/2F(0)=1-(1/e)>1/2).
Consequences of Theorem 11
Theorem 11 does not just widen the scope of our framework: it hides a crucial message for the experimenter and experimental design in a field where variations to DPO are numerous. Experiments would start by the choice that defines the loss, $\psi$ (18), from which embedding it in our framework would necessitate to find a working model of social choice whose choice probabilities have a KLST structure. Given a reward / utility function $u$ (implicit), Theorem 11 says that this second step is not necessary: any $\psi$ applicable according to Theorem 11 can operate with any model whose choice probabilities have a KLST structure, the hidden "link" operating the canonical or composite connection being a Savage loss.
4 Discussion
In this section, we discuss several points related to our KLST and Savage (KLST+S) generalization of DPO, starting with two sections on how to make the most of the generalization, i.e. craft any *PO approach that would comply with KLST+S. The two first ones, pmpo and pppo, summarize how to make the most of the setting using a composite or a proper connection (Sections 4.1, 4.2). The second one shows how to make the most of the setting using a simple knob, $\phi$-po. One section then follows on how KLST+S specializes to different approaches already published and a last section discussed pitfalls of modelings that fall outside KLST+S.
          pmpo     pppo     $\phi$-po           Ingredients     $\psi\nearrow\nearrow$     $(\ell^{a}_{0},\ell^{a}_{1})$ strictly proper     $\phi$ eligible             $\bm{\ell}^{b}$ proper     Def. 14           $\psi$ in (18)     $\psi$     $\phi^{\star}$ (Thm. 7)     Thm. 15             ($\ell^{a}_{0}$, $F$: Thm. 6)     ($F=(\ell^{a}_{0}(p)-\ell^{a}_{1}(p))^{-1}$)     & Def. 17           $G_{i}$ in (18)     $\ell^{b}_{i}$     Def. 17
Figure 2: Summary of all recipes in Sections 4.1, 4.2 and 4.3, with "ingredients" and how to define the key parts of loss $I(\pi_{\bm{\theta}})$ in (18). See text for details.
4.1 Proper-Monotone Preference Optimization (pmpo)
This is the generalization of DPO which relies on composite connections (Section 3.5). It is the most general from both Savage’s and KLST+S’s standpoints, which covers all of properness and all of the relevant functions $F$ in Theorem 6.
Definition 12.
For any strictly increasing $\psi:\mathbb{R}\rightarrow\mathbb{R}$ and $\bm{\ell}^{b}$ (proper), the pmpo Preference Optimization framework is the setting of the following parameters: $\bm{\ell}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\bm{\ell}^{b}$ in Theorem 5 and $\psi$ used as is in (18).
In this most general framework, we do not need to specify $(\ell^{a}_{0},\ell^{a}_{1})$ or $F$ since they are "hidden" in $\psi$. If necessary, this can however easily be achieved, by picking any strictly increasing function $F:\mathbb{R}\rightarrow[0,1]$ satisfying $\mathrm{dom}F\subseteq\mathrm{dom}\psi$ and $F(-z)+F(z)\leq 1,\forall z\in\mathrm{dom}F$ (Theorem 11) and building $(\ell^{a}_{0},\ell^{a}_{1})$ following Theorem 11 (the construction is given in the Appendix, Lemma Y).
4.2 Proper-Proper Preference Optimization (pppo)
This is the generalization of DPO in the canonical connection (Section 3.4). It is the most general from Savage’s standpoint in the context of KLST+S. We only need to satisfy the choice of Corollary 8 that ties Steps (2) and (3).
Definition 13.
For any losses $(\ell^{a}_{0},\ell^{a}_{1})$ (strictly proper) and $\bm{\ell}^{b}$ (proper), the pppo Preference Optimization framework is the setting of the following parameters in Theorems 5, 6 and 7: (i) $\bm{\ell}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\bm{\ell}^{b}$ in Theorem 5, (ii) $F\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\ell^{a}_{0}(p)-\ell^{a}_{1}(p))^{-1}$ in Theorem 6, (iii) $(\ell_{0},\ell_{1})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\ell^{a}_{0},\ell^{a}_{1})$ in Theorem 7.
Examples of existing techniques that fall into pppo are given in Subsection 4.4.
4.3 $\phi$-Preference Optimization ($\phi$-po)
pmpo  and pppo require to find, tune or learn two proper losses. Properness is not necessarily easy to check or ensure at training time, in particular for methods that directly learn (parts of) a loss (Kakade et al., 2011; Nock and Menon, 2020). So, we design here a simple and general knob which allows to then broadly deliver on the full framework of KLST+S preference optimization.
Definition 14.
A function $\phi$ is called eligible iff it satisfies: (i) $\mathrm{dom}\phi\supseteq[0,1]$, (ii) it is strictly convex and (iii) $\exists H\in\partial\phi$ invertible in closed form and such that $H(1/2+\delta)+H(1/2-\delta)\geq 0,\forall\delta\in(0,1/2)$.
Theorem 15.
For any eligible $\phi$, the following loss $(\ell_{0},\ell_{1})$ is strictly proper
$\ell_{1}(p)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}-\phi(p)-(1-p)H(p)$ , $\ell_{0}(p)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}-\phi(p)+pH(p),$ (20)
for any $H\in\partial\phi$. Note that if $|\phi(0)|\ll\infty,|\phi(1)|\ll\infty$, then we have the relationship $\ell_{v}(p)=D_{\phi}(v\|p)-\phi(v),\quad\forall v\in\{0,1\}$. Finally, if we add to eligibility the constraint (iv) $\phi$ is symmetric around $x=1/2$, then the loss as defined is symmetric.
Proof in Appendix, Section VII.5. Note that there are several results of close nature to that of Theorem 15, since Savage in fact, but none has the exact shape we needed here (Reid and Williamson, 2011, Theorem 7), (Williamson et al., 2016, Proposition 7), (Savage, 1971). This Theorem is instrumental in getting the proof of Theorem 11.
Remark 16.
Conditions (iii) is necessary for our specific context but the Theorem remains true even if we scrap it, as well as strictness: in essence, any convex function can be used to create a proper loss. Hence, proper "compositeness" (Reid and Williamson, 2010), which adds another function to the design of proper losses, is really only useful for technical convenience in loss design – or to correct loss defects (Bao and Charoenphakdee, 2025).
Definition 17.
The $\phi$-po Preference Optimization framework is pppo with $(\ell^{a}_{0},\ell^{a}_{1})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\ell_{0},\ell_{1})$ as per (20) and $\bm{\ell}^{b}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}[\ell^{b}_{i}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\tilde{\ell}_{1}]_{i=1}^{n}$, where $\tilde{\ell}_{1}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(-\phi(p)-\phi(1-p)-(1-p)(H(p)-H(1-p)))/2$.
Note that $\tilde{\ell}_{1}$ symmetrizes the loss as remarked in Theorem 15. When $\phi$ is differentiable, we immediately get $\phi^{\star}(z)=z\phi^{\prime-1}(z)-\phi(\phi^{\prime-1}(z))$ for the convex conjugate. Under additional assumptions on $\phi$, functions $G_{i}$ in Theorem 5 adopt a remarkable analytical form.
Corollary 18.
With the choices of Definition 17, if $\phi$ is symmetric and twice differentiable, then functions $G_{i}$ in Theorem 5 simplify as $G_{i}(p)=\phi(p)+(1-p)\phi^{\prime}(p)+p(1-p)\phi^{\prime\prime}(p),\forall i=1,2,...,n$.
Note also that the loss $\bm{\ell}^{b}$, though simple in design, is not separable if $n>2$ and amounts to designing a loss for $n$ 1-vs-the-rest classification problems. Such a design trick is used in (Hong et al., 2024).
4.4 The DPO menagerie in the context of KLST+S
We relate followers of DPO, and techniques built on top of DPO, in the context of the KLST+S framework. Other techniques, that represent an orthogonal direction, such as adding a layer to optimize marginals or posteriors on states (Ramesh et al., 2024; Zhou et al., 2024), or involving a simple autoregressive decomposition of the probability of actions (Lu et al., 2024), or replacing the reference $\pi_{\mbox{\tiny{ref}}}$ by an ad-hoc choice (such as the uniform distribution, (Lee et al., 2025)), or framing time varying preferences on top of a stochastic choice model (Son et al., 2024) can be trivially branched on KLST+S. We start by two techniques built on top of DPO, introducing home advantages or correction for length (note that some papers do both, (Lee et al., 2025)).
Home advantages / margins via properness Home advantages have been introduced by Meng et al. (2024), whereby function ${\color[rgb]{0,0,1}\psi}(z)$ in (10) includes a slack in its argument, ${\color[rgb]{0,0,1}\psi}(z-\gamma)$, transposing to DPO the ad-hoc "home advantage" introduced in data analysis (Agresti, 2012, pp 438), with $\gamma$ a so-called margin term. There is no formal justification for those, but it turns out that this "slack" is authorized in the KLST+S setting. We begin by a trivial Lemma (proof omitted).
Lemma 19.
Let $(\ell_{0},\ell_{1})$ be proper (resp. strictly proper), and $a>0$ and $c\in\mathbb{R}$ any constants. Then the loss $(\tilde{\ell}_{0},\tilde{\ell}_{1})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(a\ell_{0},a\ell_{1}+c)$ is proper (resp. strictly proper) and furthermore, using notations from Theorem 7 we have the following key quantities for loss $(\tilde{\ell}_{0},\tilde{\ell}_{1})$: $\tilde{H}^{-1}(z)=H^{-1}((z+c)/a)$ and $\tilde{\phi}^{\star}(z)=a\phi^{\star}((z+c)/a)$.
Hence the loss setting of (Meng et al., 2024) falls into pppo as long as $c/a$, which is the $\gamma$ term in SimPO, is non-negative for the stochastic choice model to be KLST  compliant. This is the case in all the papers we have spotted using the margin term (Meng et al., 2024; Zhao et al., 2025). This can also support to some extent techniques that would tune the slack (Cui et al., 2024), i.e. replacing constants $a,c$ by functions.
Length normalization via Step (1) Length normalization has been introduced in several ways  [^3] The usual ad-hoc justification is reward modification (Park et al., 2024). to express the fact that $y_{i}$ is the "level" at which preferences are formulated but $\pi(y_{i}|.)$ is a compound probability aggregating those of "elementary" actions $y_{i,k},k=1,2,...,n_{i}$ — tokens in the jargon of LLMs (Meng et al., 2024; Gu et al., 2024; Park et al., 2024; Li et al., 2025). No such dependence is reflected in any of the three steps in Section 3. We show that a simple reuse of Step (1) leads to popular corrections in (Meng et al., 2024; Gu et al., 2024; Park et al., 2024; Li et al., 2025). We first observe that for autoregressive models, we have the dependence
$\pi(y_{i}|x)=\prod_{k=1}^{n_{i}}\pi(y_{i,k}|y_{i,<k},x),\forall i=1,2,...,n_{i}.$ (21)
Suppose we wish to compute some $\tilde{\pi}(y_{i}|x)$ that would satisfy two constraints: (a) it is the best probability approximation to all $\{\pi(y_{i,k}|y_{i,<k},x)\}_{i=1}^{n_{i}}$, $\tilde{\pi}(y_{i}|x)$ and (b) it has the form $\tilde{\pi}(y_{i}|x)=u(\pi(y_{i}|x))$ via (21), for some $u$ to specify that would carry some knowledge about $\{\pi(y_{i,k}|y_{i,<k},x)\}_{i=1}^{n_{i}}$ – the simplest of which being the size, $n_{i}$. Plugging $\tilde{\pi}(y_{i}|x)=u(\pi(y_{i}|x))$ in (16) would then carry this knowledge at zero cost in the DPO pipeline. Formulating (a) is just a matter of reusing (11) by using $\pi(y_{..}|y_{..},x)$ as reference policy $\pi_{\mbox{\tiny{ref}}}$, assuming all rewards equal — and of course after Theorem 5, using a Bregman divergence $D_{\phi,\bm{G}}$ as criterion so we end up looking at the argument probability maximizing (for all i=1,2,,ni=1,2,...,n)
$J_{\phi}(\tilde{\pi}(y_{i}|x))\stackrel{{\scriptstyle\mathrm{.}}}{{=}}-\sum_{k=1}^{n_{i}}D_{\phi,\bm{G}}(\tilde{\pi}(y_{i}|x)\|\pi(y_{i,k}|y_{i,<k},x)).$ (22)
We show that the most popular corrections happen to be solutions for the KL divergence (Meng et al., 2024) and the Itakura-Saito (IS) divergence (Park et al., 2024). Plus, an advantage of using the Bregman formulation is that we do not need Lagrange multipliers to enforce a solution in $[0,1]$.
Lemma 20.
Take $\varphi(\bm{p})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\sum_{i}p_{i}\log p_{i}$. Then the probability maximizing (22) satisfies $\tilde{\pi}(y_{i}|x)=\pi(y_{i}|x)^{1/n_{i}}$. If $\varphi(\bm{p})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\sum_{i}-\log p_{i}$ and let $\alpha_{i}$ be such that $(1/n_{i})\sum_{l=1}^{n_{i}}\prod_{k\neq l}\pi(y_{i,k}|y_{i,<k},x)=\exp(-\alpha_{i}n_{i})$. Then the probability maximizing (22) satisfies $\tilde{\pi}(y_{i}|x)=\pi(y_{i}|x)\cdot\exp(\alpha_{i}n_{i})$.
(Proof in Appendix, Section VII.6) Hence $\log\tilde{\pi}(y_{i}|x)=(1/n_{i})\log\pi(y_{i}|x)$ for the KL divergence, which is the approach of Meng et al. (2024), and $\log\tilde{\pi}(y_{i}|x)=\log\pi(y_{i}|x)+\alpha_{i}n_{i}$ for the IS divergence, which is the approach of Park et al. (2024); Chen et al. (2025). Another advantage of a principled approach is that it can help to understand "default" ranges for hyperparameters: $\alpha_{i}\geq 0$ in Lemma 20, which is the approach chosen by Park et al. (2024), but not by Chen et al. (2025). The geometric average being zero-attracting, this suggests large possible ranges (if $\pi(y_{i,k}|y_{i,<k},x)=v,\forall k$, then $\alpha_{i}=\log(1/v)$).
      Loss     $\ell_{i}(\bm{p})$ proper if     $F(z)$     $\phi^{\star}(-z)$
        $n=2$ ? $n>2$ ?
      log     $\log(p_{i})$     $1/(1+\exp(-z))$     $\log(1+\exp(-z))$     (Rafailov et al., 2023)
    binary entropy     $\log(p_{i})+\sum_{j\neq i}-\log\left(1-p_{j}\right)$     $1/(1+\exp(-z/2))$     $2\log(1+\exp(-z/2))$     (Hong et al., 2024)
    square ($\tau>0$)     $\frac{1}{\tau}\cdot(1-p_{i})^{2}$     $\max\left\{0,\min\left\{\frac{1+\tau z}{2},1\right\}\right\}$     $\tau\cdot\left\{\begin{array}[]{ccl}\frac{1}{4\tau}-\frac{{x}}{\tau}&\mbox{ if }&{x}<0\\ \left({x}-\frac{1}{2\tau}\right)^{2}&\mbox{ if }&0\leq{x}\leq\frac{1}{2\tau}\\ 0&\mbox{ if }&{x}>\frac{1}{2\tau}\end{array}\right$.     (Azar et al., 2024)
    Matushita ($\mu\geq 0$)     $\mu\cdot\sqrt{\frac{1-p_{i}}{p_{i}}}$     $\frac{1}{2}\cdot\left(1+\frac{x}{\sqrt{x^{2}+\mu^{2}}}\right)$     $\frac{-z+\sqrt{\mu^{2}+z^{2}}}{2}$     (Yuan et al., 2023)
    alpha ($\beta\geq 0$)     $(1-p_{i}^{-\beta})/\beta$ / /     No general closed form     (Gupta et al., 2025)
Table 1: Examples of losses used in various variants of DPO (one example citation is given for each). Each loss may be used either as $\bm{\ell}^{a}$ or $\bm{\ell}^{b}$ in pppo (Section 4.2). Parameterized losses may require passing their parameter to the limit (see text for details). The properness "regime" is indicated for each loss, either for $n=2$ (typically for $(\ell^{a}_{0},\ell^{a}_{1})$ in pppo) or $n>2$ (for $\bm{\ell}^{b}$ in pppo).
A brief taxonomy of DPO variants Table 1 displays that many examples of DPO variations can be connected to our pppo setting, but they are not "far" from DPO (typically keeping one of its losses used in $\bm{\ell}^{a}$ and $\bm{\ell}^{b}$). The log loss is used for $\bm{\ell}^{a}$ and $\bm{\ell}^{b}$ in DPO Rafailov et al. (2023) (and variants like (Liu et al., 2025)). RRHF keeps $\bm{\ell}^{b}$ from DPO but uses Matushita for $\bm{\ell}^{a}$ ($\lim_{\mu\rightarrow 0}\phi^{\star}(-z)=\max\{0,-z\}$). ORPO (Hong et al., 2024) is an interesting case which keeps $\bm{\ell}^{a}$ from DPO but uses the binary entropy for $\bm{\ell}^{b}$. IPO (Azar et al., 2024) is also interesting because, while it keeps $\bm{\ell}^{b}$ from DPO but picks the square loss for $\bm{\ell}^{a}$, scraps the asymptotes from the convex surrogate. Finally, AlphaPO is also interesting as it keeps DPO’s $\bm{\ell}^{a}$ but replaces the log-loss in $\bm{\ell}^{b}$ by a tempered variant (Nock et al., 2023) which was introduced in Arimoto (1971, eq. (2.10)). We comment below on the "interestingness" of DPO’s variants.
4.5 Cautionary places and workarounds in the context of KLST+S
We now present a few examples of places in the KLST+S framework where one would have to be cautious about related choices.
Proper vs separable DPO (and RLHF to a large extent (Ouyang et al., 2022)) has somehow implicitly established the separability of the divergence used for its key trick in Step (1) (loss $\bm{\ell}^{b}$ in pppo). The issue with this stems from the following Theorem, which has been known and shown under various restrictive flavors for almost seven decades.
Theorem 21.
For any loss $L$ expressed as in (1), the following holds true:
  • (I)
    suppose $\ell_{i}(\bm{q})$ depends only on $q_{i}$ in (1) ($L$ is separable) and $L$ is $\bm{p}$-proper for some target vector $\bm{p}\in\Delta_{n}$ with non-zero coordinates at indexes $\mathbb{I}\subseteq\{1,2,...,n\}$. If $\mathrm{Card}(\mathbb{I})>2$, then $\ell_{i}(z)=-K_{1}\log(z)+K_{2},\forall i\in\mathbb{I}$, where $K_{2}\in\mathbb{R}$ and $K_{1}>0$ are constants;
  • (II)
    if $L$ is proper, it can be represented as $\bm{\ell}(\bm{q})=U(\bm{q})\cdot\bm{1}+[u_{i}(\bm{q})]_{i}$, such that its ABF representation $D_{\phi}(\bm{p}\|\bm{q})$ satisfies $\phi=-U(\bm{p})+\bm{p}^{\top}(-\bm{u}(\bm{p}))$, $\varphi^{\star}(-\bm{u}(\bm{q}))=U(\bm{q})$ and $\bm{u}\in\partial\phi$.
(Proof in Appendix, Section VII.7, including remarks on previous versions) Theorem 21 is important because it shows that departing from KL imposes departing from the DPO’s blueprint for losses. This has little to zero impact on the papers essentially keeping the KL divergence for $\bm{\ell}^{b}$ or DPO’s log-ratio policy dependence (Azar et al., 2024; Chen et al., 2024; Sun et al., 2025; Ethayarajh et al., 2024; Meng et al., 2024; Zhao et al., 2023; Xu et al., 2024a; Zhao et al., 2025; Slocum et al., 2025; Shao et al., 2025; Choi et al., 2025; Yang et al., 2025; Xiao et al., 2025). However, for others clearly wanting to depart from KL, the risk is real to break properness (Alfano et al., 2025; Gupta et al., 2025). Separability has a "DPO-compliant" computational advantage: the final loss’ pointwise value does not depend on unseen actions (17). Fortunately, there is a simple way to depart from KL, keep properness for a huge set of losses and keep this advantage. It exploits a simple trick well known in supervised learning.
Lemma 22.
Define $\bm{\ell}^{b}$ in Definition 13 from any loss $(\ell_{0},\ell_{1})$ as $\ell^{b}_{i}(\bm{p})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{0}(p_{i})+\sum_{j\neq i}\ell_{1}\left(1-p_{j}\right)$. If $(\ell_{0},\ell_{1})$ is proper (resp. strictly proper), then $\bm{\ell}^{b}$ is proper (resp. strictly proper).
(Proof in Appendix, Section VII.8) The proof, straightforward, shows $L(\bm{p},\bm{p})=\sum_{i}p_{i}\ell_{0}(p_{i})+(1-p_{i})\ell_{1}\left(1-p_{i}\right)$, which indeed keeps the advantage set above. ORPO uses this design (Table 1, (Hong et al., 2024)).
Proper vs improper vs twist proper Switching the standard choices of DPO can yield interesting variants, in particular when parameterizing the loss: AlphaPO (Gupta et al., 2025) replaces the log loss in $\bm{\ell}^{b}$ by a tempered version which was introduced in Arimoto (1971, eq. (2.10)), the alpha-loss (Table 1). The reason why we put the " /" is because the alpha-loss is proper only for the DPO choice (Sypherd et al., 2022). Otherwise, it is improper and the best possible $\bm{q}$ according to (1) is not the target $\bm{p}$ and it can be arbitrarily different depending on $\beta$. This property, which is of course undesirable in our setting, gives the alpha-loss a very desirable property if it is known that the preference data is subject to twists ("noise", drift, mistakes, editions / hacking, unfairness, etc.). If $\beta$ can be learned, then the loss can provably correct any twist (Soen et al., 2022), a property known as twist properness (Sypherd et al., 2022). Papers have started to imagine corrections in such settings (Kim et al., 2025) (who clamp the truth $\bm{p}$ in (1), with the change $p_{i}\rightarrow 1_{p\geq\tau}\cdot p_{i}+1_{p<\tau}\cdot(1-\alpha)$, $1_{.}$ being the indicator variable), but it can be shown that it is not twist proper.
Simplifications in the properness regime Some papers approximate functions compared to their proper "self". Such is the case for IPO (Azar et al., 2024), which discards the asymptotes of the convex conjugate (Table 1). Depending on the argument (and therefore how $\tau$ is chosen), it is of course possible that optimization stays in the properness regime. However, stepping out of the properness regime may create misbehaviors for optimization that have been long identified in supervised learning (Hastie et al., 2002, pp 348).
5 Limitations, conclusion and open problems
DPO has has a substantial impact on training LLMs (Tunstall et al., 2024; Ivison et al., 2023; Grattafiori et al., 2024; Nvidia et al., 2024; Xu et al., 2025), and an even broader impact on research, with countless variants. It is thus of primary importance to define a wide "operating" normative framework in which any such variant may be framed and understood.
We show that DPO is a particular point in a huge panorama of losses and preference models, that everywhere keeps all of DPO’s neat design tricks and properties. Why does this matter ? Zooming out from DPO to the level of generality of this panorama, it is striking that most techniques that have built over DPO look in fine extremely close to DPO (for example, they all rely on convex objectives while our framework accomodates for non-convexity). There is thus a large unexplored map to explore for substantial variations to the specific DPO choices, and this map can accommodate a number of additional requirements that have been formulated in the realm of LLMs (it naturally embeds home advantages and length correction; it is not affected by "orthogonal" extensions to DPO like marginal/posterior optimization, noise handling, time-varying preferences, etc., that can be used without change). Given the variety of environments, requirements, data specificities of LLM training, development, and deployment, it is important to frame the full freedom offered, to then be able to make choices without losing any of its advantages. Such freedom is mandatory to be able escape eventual pitfalls of specific choices (Huang et al., 2025; Alfano et al., 2025; Gupta et al., 2025), and it does indeed offer concrete leverage (Mitchell, 2023).
In terms of social impact, it would arguably be "distant" to the fundamentals we develop, even when one could expect in fine positive impacts to better understanding when preference optimization techniques can work (or not — and why), and how to design new techniques that fit in a principled framework with formal guarantees. This being said, our approach is kept it in the "straight jacket" imposed by the normative theories we use (Savage, 1971; Doignon and Falmagne, 1974; Krantz et al., 1989; Machina, 1985) – however general they arguably may be. Further, it would be quite impossible to extend our results using Savage’s theory. What this means is that some DPO variants definitely escape our framework (Huang et al., 2025; Wang et al., 2024). It does not mean that the quality of their experimental results cannot be obtained from some technique that would belong to the panorama we cover, it also does not mean that those techniques are without any connection to our framework [^4] For example, since $\log(z)=\log(z-1+1)\sim_{1}(z-1)-(z-1)^{2}/2+o((z-1)^{2})$, we easily derive that the $\chi^{2}$ divergence is a second-order approximation of a KL divergence, especially good near its optimum Huang et al. (2025).. It just means that changes must be done to our setting to accommodate those in a normative framework, which is definitely an interesting open problem.
Acknowledgments
The authors warmly thank Robert C. Williamson for many comments and discussions about this material.
Appendix
To differentiate with the numberings in the main file, the numbering of Theorems, etc. is letter-based (A, B, …).
Supplementary material on KLST  Pg VI
Supplementary material on proofs  Pg VII
$\hookrightarrow$ Proof of Theorem 5  Pg VII.1
$\hookrightarrow$ Proof of Theorem 6  Pg VII.2
$\hookrightarrow$ Proof of Theorem 7  Pg VII.3
$\hookrightarrow$ Proof of Theorem 11  Pg VII.4
$\hookrightarrow$ Proof of Theorem 15  Pg VII.5
$\hookrightarrow$ Proof of Lemma 20  Pg VII.6
$\hookrightarrow$ Proof of Theorem 21  Pg VII.7
$\hookrightarrow$ Proof of Lemma 22  Pg VII.8
VI Supplementary material on KLST
Figure 3: The monotonicity property implicitly creates a real valuation of the "niceness" of alternatives in $\mathcal{Y}^{\alpha}$. For example, computing $p({\color[rgb]{1,0,0}L_{1}}\succ{\color[rgb]{1,.5,0}L_{2}}|x)$ amounts to making a difference between the mappings of $\color[rgb]{1,0,0}L_{1}$} and $\color[rgb]{1,.5,0}L_{2}$} (left, in red). The inequality in (iii), shown in the red rectangle, establishes an order between the related differences along the axis (left), and similarly for the blue rectangle. The mapping then authorizes to compare new related differences, hence probabilities, and derive the right proposition in (7) (main file).
Figure 3 shows what monotonicity creates "behind the curtain" a mapping for the elements in $\mathcal{Y}^{\alpha}$ on a real axis, representing their "niceness" and allowing for a simple comparison of choice probabilities. We have adopted the same polarity for the difference of the valuations (all arrows go up), but the picture would lead to the same conclusion had we flipped some of them.
VII Supplementary material on proofs
VII.1 Proof of Theorem 5
Williamson et al. (2016, Proposition 7) shows the standard connection, but with our formalism, we need to also show that $\bm{\ell}\in\partial-L$, which is simple but yet needs to be shown for completeness. Properness condition imposes $\bm{p}^{\top}\bm{\ell}(\bm{p})\leq\bm{p}^{\top}\bm{\ell}(\bm{q})$ for all $\bm{p},\bm{q}\in\Delta_{n}$. Put all on the right hand side and add there $L(\bm{q},\bm{q})+L(\bm{q},\bm{q})$. Expand with the definition of $L$ and it yields:
$-\bm{p}^{\top}\bm{\ell}(\bm{p})-(-\bm{q}^{\top}\bm{\ell}(\bm{q}))-(\bm{p}-\bm{q})^{\top}\bm{\ell}(\bm{q})$ $\geq$ $0,$
which implies indeed $\bm{\ell}\in\partial-L$. The convexity of $L(\bm{q},\bm{q})$ was originally shown in Gneiting and Raftery (2007).
VII.2 Proof of Theorem 6
We exploit a classical result of Krantz et al. (1989, Chapter 17, Theorem 2), in two steps: we first that the local choice structure and monotonicity of lotteries extends to alternatives of $\mathcal{Y}$ (i.e. for $\alpha=\{0,1\}$) and then we show that considering lotteries allows to bypass the use of a quite non intuitive solvability axiom (we separately note that it also appears in Debreu (1958)’s theory, thus when zero abstention is enforced).
Let us start by step one. We cover the properties involved in the local choice structure and monotonicity for lotteries and show that they specialize to the lotteries atomic components, i.e. they hold for $\mathcal{Y}$ as well. The following simple Lemma is central to our results. It actually contains more results than strictly needed for our purpose as those shall be used in a discussion comparing the original KLST model to our lottery model. One can infer from the definition of zero abstention that it is a stronger notion than preference (zero abstention implies a preference from its definition (5), but the converse is not necessarily true). The Lemma shows such a property holds at the "deeper” stage of alternatives.
Lemma W.
$\forall{\color[rgb]{1,0,0}y_{1}},{\color[rgb]{1,.5,0}y_{2}},{\color[rgb]{1.0,0.88,0.21}y_{3}},{\color[rgb]{0,0.8046875,0}y_{4}}\in\mathcal{Y}$, the following holds
  • (I)
    on preferences (P):
    $\forall\alpha\in(0,1):({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}\triangleright_{x}({\color[rgb]{1.0,0.88,0.21}y_{3}}{\color[rgb]{0,0.8046875,0}y_{4}})_{\alpha}$ $\Rightarrow$ $({\color[rgb]{1,0,0}y_{1}}\triangleright_{x}{\color[rgb]{1.0,0.88,0.21}y_{3}})\wedge({\color[rgb]{1,.5,0}y_{2}}\triangleright_{x}{\color[rgb]{0,0.8046875,0}y_{4}}),$ (23)
    $\exists\alpha\in(0,1):({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}\triangleright_{x}({\color[rgb]{1.0,0.88,0.21}y_{3}}{\color[rgb]{0,0.8046875,0}y_{4}})_{\alpha}$ $\Leftrightarrow$ $({\color[rgb]{1,0,0}y_{1}}\triangleright_{x}{\color[rgb]{1.0,0.88,0.21}y_{3}})\vee({\color[rgb]{1,0,0}y_{1}}\triangleright_{x}{\color[rgb]{0,0.8046875,0}y_{4}})\vee({\color[rgb]{1,.5,0}y_{2}}\triangleright_{x}{\color[rgb]{1.0,0.88,0.21}y_{3}})\vee({\color[rgb]{1,.5,0}y_{2}}\triangleright_{x}{\color[rgb]{0,0.8046875,0}y_{4}}).$ (24)
    Furthermore, if among the four alternatives / actions in (24), three are identical and Bearability holds, then the equivalence holds with $\triangleright_{x}$ between the two distinct elements on the right proposition: for example, if $y\stackrel{{\scriptstyle\mathrm{.}}}{{=}}{\color[rgb]{1,0,0}y_{1}}={\color[rgb]{1,.5,0}y_{2}}={\color[rgb]{1.0,0.88,0.21}y_{3}}\neq{\color[rgb]{0,0.8046875,0}y_{4}}$, then the RHS is $y\triangleright_{x}{\color[rgb]{0,0.8046875,0}y_{4}}$;
  • (II)
    on zero abstention (ZA):
    $\exists\alpha\in(0,1):(({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha},({\color[rgb]{1.0,0.88,0.21}y_{3}}{\color[rgb]{0,0.8046875,0}y_{4}})_{\alpha})\in\mathcal{E}_{x}^{\alpha}$ $\Leftrightarrow$ $(({\color[rgb]{1,0,0}y_{1}},{\color[rgb]{1.0,0.88,0.21}y_{3}})\in\mathcal{E}_{x})\wedge(({\color[rgb]{1,0,0}y_{1}},{\color[rgb]{0,0.8046875,0}y_{4}})\in\mathcal{E}_{x})$ (25)
    $\wedge(({\color[rgb]{1,.5,0}y_{2}},{\color[rgb]{1.0,0.88,0.21}y_{3}})\in\mathcal{E}_{x})\wedge(({\color[rgb]{1,.5,0}y_{2}},{\color[rgb]{0,0.8046875,0}y_{4}})\in\mathcal{E}_{x}).$
Proof.
To prove (23), we decompose using (4)
$p(({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}\succ({\color[rgb]{1.0,0.88,0.21}y_{3}}{\color[rgb]{0,0.8046875,0}y_{4}})_{\alpha}|x)$ $=$ $\alpha^{2}p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{1.0,0.88,0.21}y_{3}}|x)+\alpha(1-\alpha)p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{0,0.8046875,0}y_{4}}|x)$ (26)
$+\alpha(1-\alpha)p({\color[rgb]{1,.5,0}y_{2}}\succ{\color[rgb]{1.0,0.88,0.21}y_{3}}|x)+(1-\alpha)^{2}p({\color[rgb]{1,.5,0}y_{2}}\succ{\color[rgb]{0,0.8046875,0}y_{4}}|x),$
and we have $p(({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}\succ({\color[rgb]{1.0,0.88,0.21}y_{3}}{\color[rgb]{0,0.8046875,0}y_{4}})_{\alpha}|x)\geq 1/2$. If $p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{1.0,0.88,0.21}y_{3}}|x)<1/2$, then there are values of $\alpha$ close to 1 such that $p(({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}\succ({\color[rgb]{1.0,0.88,0.21}y_{3}}{\color[rgb]{0,0.8046875,0}y_{4}})_{\alpha}|x)<1/2$, and that is the same if $p({\color[rgb]{1,.5,0}y_{2}}\succ{\color[rgb]{0,0.8046875,0}y_{4}}|x)<1/2$ for values of $\alpha$ close to 0. Hence, we must have ${\color[rgb]{1,0,0}y_{1}}\triangleright_{x}{\color[rgb]{1.0,0.88,0.21}y_{3}}$ and ${\color[rgb]{1,.5,0}y_{2}}\triangleright_{x}{\color[rgb]{0,0.8046875,0}y_{4}}$. This is the most we can get because if $p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{1.0,0.88,0.21}y_{3}}|x)=p({\color[rgb]{1,.5,0}y_{2}}\succ{\color[rgb]{0,0.8046875,0}y_{4}}|x)=1$, then we already have $p(({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}\succ({\color[rgb]{1.0,0.88,0.21}y_{3}}{\color[rgb]{0,0.8046875,0}y_{4}})_{\alpha}|x)\geq\alpha^{2}+(1-\alpha)^{2}\geq 1/2$ so we cannot enforce either ${\color[rgb]{1,0,0}y_{1}}\triangleright_{x}{\color[rgb]{0,0.8046875,0}y_{4}}$ or ${\color[rgb]{1,.5,0}y_{2}}\triangleright_{x}{\color[rgb]{1.0,0.88,0.21}y_{3}}$.
To prove (24), we just have to remark that $\alpha^{2}+2\alpha(1-\alpha)+(1-\alpha)^{2}=1$. We then prove (25) and note that $1=p(({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}\succ({\color[rgb]{1.0,0.88,0.21}y_{3}}{\color[rgb]{0,0.8046875,0}y_{4}})_{\alpha}|x)+p(({\color[rgb]{1.0,0.88,0.21}y_{3}}{\color[rgb]{0,0.8046875,0}y_{4}})_{\alpha}\succ({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}|x)$ is equivalent to, after decomposition using (4),
$1$ $=$ $\alpha^{2}(p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{1.0,0.88,0.21}y_{3}}|x)+p({\color[rgb]{1.0,0.88,0.21}y_{3}}\succ{\color[rgb]{1,0,0}y_{1}}|x))+\alpha(1-\alpha)(p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{0,0.8046875,0}y_{4}}|x)+p({\color[rgb]{0,0.8046875,0}y_{4}}\succ{\color[rgb]{1,0,0}y_{1}}|x))$
$+\alpha(1-\alpha)(p({\color[rgb]{1,.5,0}y_{2}}\succ{\color[rgb]{1.0,0.88,0.21}y_{3}}|x)+p({\color[rgb]{1.0,0.88,0.21}y_{3}}\succ{\color[rgb]{1,.5,0}y_{2}}|x))+(1-\alpha)^{2}(p({\color[rgb]{1,.5,0}y_{2}}\succ{\color[rgb]{0,0.8046875,0}y_{4}}|x)+p({\color[rgb]{0,0.8046875,0}y_{4}}\succ{\color[rgb]{1,.5,0}y_{2}}|x)),$
which, since $\alpha^{2}+2\alpha(1-\alpha)+(1-\alpha)^{2}=1$, leads to the statement. We tackle the last statement, if among the four elements in (24), three are identical. We see from (26) two probabilities are identical (say $p$) and two are $1/2$ from assumption Bearability; furthermore the coefficient of the probability $p$ is $\alpha$ and that of $1/2$ is $1-\alpha$, or we permute those, which, if we look at inequality $p(({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}\succ({\color[rgb]{1.0,0.88,0.21}y_{3}}{\color[rgb]{0,0.8046875,0}y_{4}})_{\alpha}|x)\geq 1/2$, ends up simplifying the inequality as $p\geq 1/2$. ∎
Specialization of Bearability
It follows by considering all lotteries of the form $L\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(yy)_{\alpha}$ for $y\in\mathcal{Y}$: $(L,L)\in\mathcal{E}^{\alpha}_{x}$ implies after using the property of lotteries (4) $1\stackrel{{\scriptstyle\mathrm{.}}}{{=}}p(L\succ L|x)+p(L\succ L|x)=(\alpha^{2}+2\alpha(1-\alpha)+(1-\alpha)^{2})\cdot(p(y\succ y|x)+p(y\succ y|x))=p(y\succ y|x)+p(y\succ y|x)$, hence $(y,y)\in\mathcal{E}_{x}$ and bearability holds for $G_{x}$. Hence, Bearability holds on alternatives / actions as well.
[…]
Take ${\color[rgb]{1,0,0}y_{1}},{\color[rgb]{1,.5,0}y_{2}},{\color[rgb]{1.0,0.88,0.21}y_{3}}\in\mathcal{Y}$, with (i) $\{{\color[rgb]{1,.5,0}y_{2}},{\color[rgb]{1,0,0}y_{1}},{\color[rgb]{1.0,0.88,0.21}y_{3}}\}$ a wedge in $\mathcal{E}_{x}$ and (ii) $({\color[rgb]{1,0,0}y_{1}}\triangleright_{x}{\color[rgb]{1,.5,0}y_{2}})\wedge({\color[rgb]{1,0,0}y_{1}}\triangleright_{x}{\color[rgb]{1.0,0.88,0.21}y_{3}})$. Create three lotteries for any fixed $\alpha\in(0,1)$: $L_{1}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,0,0}y_{1}})_{\alpha}$, $L_{2}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}$ and $L_{3}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1.0,0.88,0.21}y_{3}})_{\alpha}$. Point (II) in Lemma W and (i) yield immediately that $L_{2},L_{1},L_{3}$ is a wedge in $\mathcal{E}_{x}^{\alpha}$. We also remark
$p(L_{1}\succ L_{2}|x)$ $\stackrel{{\scriptstyle\mathrm{.}}}{{=}}$ $p(({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,0,0}y_{1}})_{\alpha}\succ({\color[rgb]{1,0,0}y_{1}}{\color[rgb]{1,.5,0}y_{2}})_{\alpha}|x),$
[…]
Take any $y,y^{\prime}\in\mathcal{Y}$ such that $y\triangleright_{x}y^{\prime}$. It comes $L\triangleright_{x}L^{\prime}$ for $L\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(yy)_{\alpha}$ and $L^{\prime}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(y^{\prime}y^{\prime})_{\alpha}$. We show the result by induction on $n$. If $n=1$, then we get $L\triangleright_{x}L^{\prime}$ (which we knew) and $(L,L^{\prime})\in\mathcal{E}^{\alpha}_{x}$, which implies $(y,y^{\prime})\in\mathcal{E}_{x}$ from point (II) of Lemma W, and concludes the specialization proof for $n=1$. Suppose the property true for any $n\geq 1$ and we show it for $n+1$. Consider the path ${\color[rgb]{1,0,0}((yy)_{\alpha}=L_{0},L_{1}),(L_{1},L_{2}),...,(L_{n-1},L_{n})},(L_{n},L_{n+1}=(y^{\prime}y^{\prime})_{\alpha})$ in $\mathcal{E}^{\alpha}_{x}$ and such that $L_{i}\triangleright_{x}L_{i+1},\forall i\in[n]$. Denote for short $L_{i}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(y_{i_{1}},y_{i_{2}})$ ($y_{0_{1}}=y_{0_{2}}=y$ with a slight abuse of notation). We now exploit the induction hypothesis on the red part of the path: point (I) in Lemma W yields
$(y_{i_{1}}\triangleright_{x}y_{(i+1)_{1}})\wedge(y_{i_{2}}\triangleright_{x}y_{(i+1)_{2}}),\forall i=0,1,...,n-1,$ (27)
And point (II) in Lemma W yields
$((y_{i_{1}},y_{(i+1)_{1}})\in\mathcal{E}_{x})\wedge((y_{i_{2}},y_{(i+1)_{2}})\in\mathcal{E}_{x}),\forall i=0,1,...,n-1.$ (28)
[…]
Specialization of Monotonicity
This one is even simpler: suppose monotonicity holds for some $\alpha\in(0,1)$ and for any ${\color[rgb]{1,0,0}y_{1}},{\color[rgb]{1,.5,0}y_{2}},{\color[rgb]{1.0,0.88,0.21}y_{3}},{\color[rgb]{0,0.8046875,0}y_{4}},{\color[rgb]{0,1,1}y_{5}},{\color[rgb]{.5,0,.5}y_{6}}$ in $\mathcal{Y}$, build lotteries ${\color[rgb]{1,0,0}L_{1}}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}({\color[rgb]{1,0,0}y_{1}},{\color[rgb]{1,0,0}y_{1}})_{\alpha}$, ${\color[rgb]{1,.5,0}L_{2}}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}({\color[rgb]{1,.5,0}y_{2}},{\color[rgb]{1,.5,0}y_{2}})_{\alpha}$, and so on until ${\color[rgb]{.5,0,.5}L_{6}}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}({\color[rgb]{.5,0,.5}y_{6}},{\color[rgb]{.5,0,.5}y_{6}})_{\alpha}$. Using the fact that, by definition and for example
$p({\color[rgb]{1,0,0}L_{1}}\succ{\color[rgb]{1,.5,0}L_{2}}|x)$ $=$ $(\alpha^{2}+2\alpha(1-\alpha)+(1-\alpha)^{2})\cdot p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{1,.5,0}y_{2}}|x)=p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{1,.5,0}y_{2}}|x),$
we easily get (iii) and (iv) valid over the corresponding alternatives / actions as well. Point (ii) in Lemma W allows to conclude that points (i) (triangle) and (ii) (wedge) of monotonicity also transfers on to alternatives / actions.
At this point, we have completed the first step of our proof. We go on to the second step, showing first that expandability implies the following result on $\mathcal{Y}$: suppose $p_{1}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}p({\color[rgb]{1,0,0}y_{1}}\succ y|x)<p({\color[rgb]{1,.5,0}y_{2}}\succ y|x)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}p_{2}$. Then for any $q\in(p_{1},p_{2})$, if we fix
$\alpha$ $=$ $\frac{q-p_{1}}{p_{2}-p_{1}}\quad\left(\alpha\in(0,1)\right),$ (29)
then the lottery $({\color[rgb]{1,.5,0}y_{2}}{\color[rgb]{1,0,0}y_{1}})_{\alpha}$ satisfies $p\left(({\color[rgb]{1,.5,0}y_{2}}{\color[rgb]{1,0,0}y_{1}})_{\alpha}\succ y|x\right)=q$ from (4). This shows that (Krantz et al., 1989, Chapter 17, Theorem 2)’s solvability axiom (3) is satisfied. Our definition of local choice structure implies their axioms (1) and (4) and their monotonicity axiom is ours and so their Theorem 2 implies the existence of $u(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ such that for any ${\color[rgb]{1,0,0}y_{1}},{\color[rgb]{1,.5,0}y_{2}},{\color[rgb]{0,0.8046875,0}y_{3}},{\color[rgb]{.5,0,.5}y_{4}}\in\mathcal{Y}$ and $x\in\mathcal{X}$,
$p({\color[rgb]{1,0,0}y_{1}}\succ{\color[rgb]{1,.5,0}y_{2}}|x)\leq p({\color[rgb]{0,0.8046875,0}y_{3}}\succ{\color[rgb]{.5,0,.5}y_{4}}|x)$ $\Leftrightarrow$ $u({\color[rgb]{1,0,0}y_{1}},x)-u({\color[rgb]{1,.5,0}y_{2}},x)\leq u({\color[rgb]{0,0.8046875,0}y_{3}},x)-u({\color[rgb]{.5,0,.5}y_{4}},x)$ (30)
(and furthermore equality can only happen simultaneously on both sides). Consider the 2D plot giving $p(y\succ y^{\prime}|x)$ as a function of $u(y,x)-u(y^{\prime},x)$. (30) implies that it can be intrapolated by a strictly increasing function $F$. We also have $F(u(y,x)-u(y^{\prime},x))+F(u(y^{\prime},x)-u(y,x))=p(y\succ y^{\prime}|x)+p(y^{\prime}\succ y|x)\leq 1$ so $F(z)+F(-z)\leq 1$. This ends the proof of Theorem 6.
Remark X.
[…]
VII.3 Proof of Theorem 7
Let us first show that
$\ell_{0}(p)-\ell_{1}(p)$ $\in$ $\partial\phi(p).$
This is equivalent to show that for any $q,p$,
$\phi(q)-\phi(p)-(q-p)\cdot(\ell_{0}(p)-\ell_{1}(p))$ $\geq$ $0.$
Using $\phi(v)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}-v\ell_{1}(v)-(1-v)\ell_{0}(v)$, this turns into the inequality
$-q\ell_{1}(q)-(1-q)\ell_{0}(q)+\underbrace{p\ell_{1}(p)+(1-p)\ell_{0}(p)-(q-p)\cdot(\ell_{0}(p)-\ell_{1}(p))}_{=q\ell_{1}(p)+(1-q)\ell_{0}(p)}$ $\geq$ $0,$
which becomes equivalently $q\ell_{1}(q)+(1-q)\ell_{0}(q)\leq q\ell_{1}(p)+(1-q)\ell_{0}(p)$, which is just the statement of properness for $\ell$ on $\Delta_{2}$ and thus shows $\ell_{0}(p)-\ell_{1}(p)\in\partial\phi(p)$. $\ell$ being strictly proper, $H(p)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{0}(p)-\ell_{1}(p)$ is invertible so if we postulate $x=H(p)$, then necessarily
$\ell_{0}(H^{-1}(x))-\ell_{1}(H^{-1}(x))$ $=$ $x.$
Suppose $p\neq 0$, multiply both sides by $H^{-1}(x)$ and reorganize:
$H^{-1}(x)\cdot\ell_{0}(H^{-1}(x))$ $=$ $x\cdot H^{-1}(x)+H^{-1}(x)\cdot\ell_{1}(H^{-1}(x)).$
Add $\ell_{0}(H^{-1}(x))$ to both sides and reorganize:
$\ell_{0}(H^{-1}(x))$ $=$ $x\cdot H^{-1}(x)+H^{-1}(x)\cdot\ell_{1}(H^{-1}(x))+(1-H^{-1}(x))\cdot\ell_{0}(H^{-1}(x))$ (31)
$=$ $x\cdot H^{-1}(x)-\phi(H^{-1}(x))$
(the last identity is the definition of $\phi$). Since $H\in\partial\phi(p)$, recovering $x=H(p)$ yields from the definition of the subdifferential that for any $t\in\mathrm{dom}\phi$,
$\phi(t)-\phi(p)-(t-p)H(p)$ $\geq$ $0,$
which reformulates as $t^{\prime}x-\phi(t^{\prime})\leq x\cdot H^{-1}(x)-\phi(H^{-1}(x))$ ($\forall t^{\prime}$), implying $x\cdot H^{-1}(x)-\phi(H^{-1}(x))=\phi^{\star}(x)$ for the right-hand side in (31), so we indeed get in the general case:
$\ell_{0}\circ H^{-1}(x)$ $=$ $\phi^{\star}(x).$
Now, if the loss is symmetric, then it has $\ell_{0}(p)=\ell_{1}(1-p)$, inducing the symmetry H(1p)=H(p)H(1-p)=-H(p), yielding
$\ell_{0}(1-H^{-1}(x))$ $=$ $\ell_{1}\circ H^{-1}(-x),$
and so $\ell\circ H^{-1}(x)=\phi^{\star}(-x)$ (with the convention $\ell\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell_{1}$), as claimed.
VII.4 Proof of Theorem 11
The Theorem relies on the following Lemma.
Lemma Y.
For any function $\ell:[0,1]\rightarrow\overline{\mathbb{R}}$ strictly increasing, there exists $\ell_{1}:[0,1]\rightarrow\overline{\mathbb{R}}$ such that $(\ell_{0}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell,\ell_{1})$ is strictly proper. More specifically, we can pick, for any constants $K\in\mathbb{R},a\in[0,1]$,
$(\ell_{0},\ell_{1})(p)$ $=$ $\left(\ell(p),K-\left(\int_{a}^{p}\frac{\ell(t)}{t^{2}}\mathrm{d}t+\frac{1-p}{p}\cdot\ell(p)\right)\right)$
Proof.
Our proof is indeed constructive: we build $\ell_{1}$ via (20) by building
$\ell\rightarrow\phi\rightarrow\ell_{1}$ (32)
We start by the left $\rightarrow$. Knowing $\ell$, $\phi$ in (20) is the solution of an ODE
$y-py^{\prime}$ $=$ $\ell$
($\phi=-y$), from which we get the general solution ($K$ is a constant)
$y$ $=$ $K\exp\left(-\int\frac{1}{-t}\mathrm{d}t\right)+\exp\left(-\int\frac{1}{-t}\mathrm{d}t\right)\cdot\int\frac{\ell(t)}{-t}\cdot\exp\left(\int\frac{1}{-t}\mathrm{d}t\right)\mathrm{d}t$
$=$ $K\cdot p-p\cdot\int\frac{\ell(t)}{t^{2}}\mathrm{d}t.$
and we check that ($a\in[0,1]$)
$\phi(p)$ $=$ $p\cdot\left(\int_{a}^{p}\frac{\ell(t)}{t^{2}}\mathrm{d}t-K\right)$
is indeed strictly convex  [^5] If $\ell$ is differentiable, we get the simple expression $\phi^{\prime\prime}(p)=\ell^{\prime}(p)/p>0$ because $\ell$ is strictly increasing.. We then compute
$H(p)$ $\stackrel{{\scriptstyle\mathrm{.}}}{{=}}$ $\phi^{\prime}(p)$
$=$ $\int_{a}^{p}\frac{\ell(t)}{t^{2}}\mathrm{d}t+\frac{\ell(p)}{p}-K$
and, given that $\ell_{0}(p)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell(p)$, we can complete (32) and get from (20)
$\ell_{1}(p)$ $\stackrel{{\scriptstyle\mathrm{.}}}{{=}}$ $-\phi(p)-(1-p)H(p)$
$=$ $-p\cdot\left(\int_{a}^{p}\frac{\ell(t)}{t^{2}}\mathrm{d}t-K\right)-(1-p)\cdot\left(\int_{a}^{p}\frac{\ell(t)}{t^{2}}\mathrm{d}t+\frac{\ell(p)}{p}-K\right)$
$=$ $K-\left(\int_{a}^{p}\frac{\ell(t)}{t^{2}}\mathrm{d}t+\frac{1-p}{p}\cdot\ell(p)\right).$
This ends the proof of the Lemma. ∎
Remark that if $\ell$ is differentiable then $\ell_{1}$ is also differentiable and we get
$\ell_{1}^{\prime}(p)$ $=$ $-\frac{1-p}{p}\cdot\ell^{\prime}(p),$
which in addition to proving that $\ell_{1}$ is strictly decreasing also checks the properness condition when $(\ell_{0}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\ell,\ell_{1})$ is differentiable (Reid and Williamson, 2010, Theorem 1). Now, $F:\mathbb{R}\rightarrow[0,1]$ being strictly increasing, $F^{-1}:[0,1]\rightarrow\mathbb{R}$ is also strictly increasing, so $\psi\circ F^{-1}$ is strictly increasing, and hence from Lemma Y there exists a strictly proper loss $(\ell_{0},\ell_{1})$ such that $\ell_{0}=\psi\circ F^{-1}$, or equivalently
$\psi$ $=$ $\ell_{0}\circ F,$
as claimed. To prove the Theorem, we just have to remark that $\ell\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\psi\circ F^{-1}$ is a function $0,1]\rightarrow\mathbb{R}$ and strictly increasing, being the composition of two strictly increasing functions, which concludes using Lemma Y.
VII.5 Proof of Theorem 15
We want $\phi(p)=p\cdot\ell_{1}(p)+(1-p)\cdot\ell_{0}(p)$ for some strictly proper loss $(\ell_{0},\ell_{1})$, that we assume wlog continuous on $(0,1)$ (because $\phi$ is necessarily continuous on $(0,1)$, being convex). We get the necessary relationship for $p<1$:
$\ell_{0}(p)$ $=$ $-\frac{\phi(p)+p\ell_{1}(p)}{1-p}.$ (33)
Picking any $H\in\partial\phi$, the subdifferential of $L(p,q)$ at $q$, $\partial_{q}L(p,q)$, contains the following selections, using (33) to get rid of $\ell_{0}$:
$\partial_{q}L(p,q)$ $\ni$ $pJ(q)-\frac{1-p}{(1-q)^{2}}\cdot\left((H(q)+\ell(q)+qJ(q))\cdot(1-q)+\phi(q)+q\ell(q)\right),$
where $J$ is the derivative of $\ell_{1}$ where it is differentiable, and otherwise an element of the sub/super differential if it is not (a real in the interval defined by the left and right derivatives, guaranteed to exist because $\ell_{1}$ is continuous). The strict properness condition, that $\{0\}=\partial_{q}L(p,q)|_{p=q}$, imposes thus
$pJ(p)-\frac{1-p}{(1-p)^{2}}\cdot\left((H(p)+\ell_{1}(p)+pJ(p))\cdot(1-p)+\phi(p)+p\ell_{1}(p)\right)$ $=$ $0,$ (34)
which simplifies in $\ell_{1}(p)=-\phi(p)-(1-p)H(p)=D_{\phi,H}(1\|p)-\phi(1)$ (if $\phi(1)$ is finite), as claimed. Using (33), we get
$\ell_{0}(p)$ $=$ $-\frac{\phi(p)+p\cdot(-\phi(p)-(1-p)H(p))}{1-p}$
$=$ $-\phi(p)+pH(p),$
as claimed.
VII.6 Proof of Lemma 20
The solution to (22) follows from a standard property of Bregman divergences (see e.g. (Nock and Menon, 2020, Appendix)):
$\tilde{\pi}(y_{i}|x)$ $=$ $\phi^{\prime-1}\left(\frac{1}{n_{i}}\cdot\sum_{k}\phi^{\prime}(\pi(y_{ik}|y_{i,<k},x))\right).$ (35)
Remark the absence of Lagrange multipliers: such generalized means have the property that the solution $\tilde{\pi}(y_{i}|x)$ is between the min and max values of $\tilde{\pi}(y_{ik}|y_{i,<k},x),j=1,2,$..., so we are guaranteed that $\tilde{\pi}(y_{i}|x)\in[0,1]$. The solution for the KL divergence is then immediate, the optimum being the geometric mean. For the Itakura-Saito (IS) divergence, we just have to rewrite the harmonic mean:
$\tilde{\pi}(y_{i}|x)=\frac{n}{\sum_{k}\frac{1}{\pi(y_{i,k}|y_{i,<k},x)}}=\gamma_{i}\cdot\prod_{k}\pi(y_{i,k}|y_{i,<k},x)=u(\pi(y_{i}|x))\mbox{ for }u(z)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}z/\gamma_{i},$
with $\gamma_{i}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(1/n)\cdot\sum_{l}\prod_{k\neq l}\pi(y_{i,k}|y_{i,<k},x)$ ($\in[0,1]$). It is hard to estimate the sum-product element but if we compute the $\alpha_{i}>0$ such that
$\gamma_{i}$ $=$ $\exp(-\alpha_{i}n_{i}),$
then we get $u(z)=z\cdot\exp(\alpha_{i}n_{i})$, as claimed.
VII.7 Proof of Theorem 21
Part I. With $\phi(\bm{p})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}-\bm{p}^{\top}\bm{\ell}(\bm{p})$, the properness condition in $\bm{q}$ (1) is equivalent to having, $\forall\bm{p},\bm{q}\in\Delta_{n}$,
$\phi(\bm{p})-\phi(\bm{q})-(\bm{p}-\bm{q})^{\top}\cdot(-\bm{\ell}(\bm{q}))$ $\geq$ $0.$ (36)
For any set $\{\bm{p}_{1},\bm{p}_{2},...,\bm{p}_{k}\}$ ($k>0$) and $\bm{r}\in\Delta_{k}$, if we let $\bm{q}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathbb{E}_{j\sim\bm{r}}[\bm{p}_{j}]\in\Delta_{n}$ and compute the expectation of the $k$ inequalities (36), we get
$\mathbb{E}_{j\sim\bm{r}}[\phi(\bm{p}_{j})-\phi(\bm{q})-(\bm{p}_{j}-\bm{q})^{\top}\cdot(-\bm{\ell}(\bm{q}))]=\mathbb{E}_{j\sim\bm{r}}[\phi(\bm{p}_{j})]-\phi(\mathbb{E}_{j\sim\bm{r}}[\bm{p}_{j}])$ $\geq$ $0,$
so $\phi$ has to be convex on $\Delta_{n}$. The target $\bm{p}$ has three or more non-zero probabilities in indexes $\mathbb{I}$. In this case, if we restrict $\bm{q}$ to have non-zero coordinates in the same indexes (allowed by the properness assumption), then any two coordinates of $\bm{q}$ on these indexes are allowed to take values in a non-empty 2-dimensional open subset of a vector space; this, along with convexity and (36), impose the subdifferential property $\bm{\ell}\in\partial\phi$ (without loss of generality, we suppose vectors restricted to indexes whose coordinates in $\bm{p}$ are non zero) which, since $\phi$ is separable and of the form $\phi(\bm{p})=-\bm{p}^{\top}\bm{\ell}(\bm{p})$, provides at least 3 properties:
$-\ell_{i}+K$ $\in$ $\partial-z\ell_{i}(z),\forall i\in\mathbb{I},$ (37)
where $K$ is a constant. We note that for all these indexes
$\partial-z\ell_{i}(z)$ $=$ $\{-\ell_{i}(z)-z\cdot h_{i}(z)\}_{h_{i}\in H_{i}},$ (38)
where $H_{i}$ is the set of functions taking value $\ell^{\prime}_{i}$ if it is differentiable or a real in the interval defined by the left and right derivatives if not. Hence (37) and (38) yield:
$\exists K_{1}\in\mathbb{R}:\forall i=1,2,...,n,\exists h_{i}\in H:h_{i}(z)=-K_{1}/z.$
$\ell_{i}$ being continuous (otherwise $\phi$ is not continuous, hence not convex), this implies $\ell_{i}(z)=-K_{1}\log(z)+K_{2}$ ($\forall i\in\mathbb{I}$) with constraints $K_{2}\in\mathbb{R}$ and $K_{1}>0$ (otherwise $\phi$ is not convex).
Remark Z.
[…]
Part II mostly follows from Williamson et al. (2016, Proposition 7). suppose $L(\bm{p},\bm{q})$ is proper. Write the vector of loss coordinates in $\bm{q}$ as
$\bm{\ell}(\bm{q})$ $\stackrel{{\scriptstyle\mathrm{.}}}{{=}}$ $U(\bm{q})\cdot\bm{1}+[u_{i}(\bm{q})]_{i},$ (39)
where $U(\bm{q})$ does not depend on the index and $u_{i}$ does (and so $\bm{p}^{\top}\bm{\ell}(\bm{q})=U(\bm{q})+\bm{p}^{\top}\bm{u}(\bm{q}),\forall\bm{p}\in\Delta_{n}$). Then $\bm{p}$-properness imposes $\forall\bm{q}\in\mathbb{R}^{n}$,
$0$ $\leq$ $L(\bm{p},\bm{q})-L(\bm{p},\bm{p})$ (40)
$=$ $U(\bm{q})+\bm{p}^{\top}\bm{u}(\bm{q})-U(\bm{p})-\bm{p}^{\top}\bm{u}(\bm{p})$
$=$ $(-U(\bm{p})+\bm{p}^{\top}(-\bm{u}(\bm{p})))-(-U(\bm{q})+\bm{q}^{\top}(-\bm{u}(\bm{q})))-(\bm{p}-\bm{q})^{\top}(-\bm{u}(\bm{q})).$ (41)
We did not put any constraint on the decomposition of loss coordinates but (41) gives us a very clear one: it must be done in such a way that (i) $U(\bm{q})+\bm{q}^{\top}(-\bm{u}(\bm{q}))$ is convex (see Part I) and (ii)
$-\bm{u}(\bm{q})$ $\in$ $\partial-U(\bm{q})+\bm{q}^{\top}(-\bm{u}(\bm{q}))\quad(\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\varphi(\bm{q})).$ (42)
Let us go back to (40) and rewrite it:
$U(\bm{q})+\bm{p}^{\top}\bm{u}(\bm{q})-U(\bm{p})-\bm{p}^{\top}\bm{u}(\bm{p})$ $=$ $\varphi(\bm{p})+U(\bm{q})-\bm{p}^{\top}(-\bm{u}(\bm{q})).$
We now show
$U(\bm{q})$ $=$ $\varphi^{\star}(-\bm{u}(\bm{q})),$ (43)
where $\varphi^{\star}(\bm{z})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\sup_{\bm{t}}\bm{t}^{\top}\bm{z}-\varphi(\bm{t})$ is the convex conjugate. This follows by remarking that (41) can be rewritten, switching $\bm{p}$ for $\bm{t}$ for clarity of the link we intend to make with the convex conjugate:
$\bm{t}^{\top}(-\bm{u}(\bm{q}))-\underbrace{(-U(\bm{t})+\bm{t}^{\top}(-\bm{u}(\bm{t})))}_{=\phi(\bm{t})}$ $\leq$ $\bm{q}^{\top}(-\bm{u}(\bm{q}))-\underbrace{(-U(\bm{q})+\bm{q}^{\top}(-\bm{u}(\bm{q})))}_{=\phi(\bm{q})},\forall\bm{t}.$
Hence,
$\varphi^{\star}(-\bm{u}(\bm{q}))$ $=$ $\bm{q}^{\top}(-\bm{u}(\bm{q}))-\phi(\bm{q})$
$=$ $U(\bm{q}),$
as claimed. We end up having $L(\bm{p},\bm{q})-L(\bm{p},\bm{p})=\varphi(\bm{p})+\varphi^{\star}(-\bm{u}(\bm{q}))-\bm{p}^{\top}(-\bm{u}(\bm{q}))=D_{\varphi}(\bm{p}\|\bm{q})$, the ABF representation of a Bregman divergence.
VII.8 Proof of Lemma 22
Showing the Lemma is a matter of writing
$L(\bm{p},\bm{q})$ $=$ $\sum_{i}p_{i}\ell_{0}(q_{i})+p_{i}\cdot\sum_{j\neq i}\ell_{1}\left(1-q_{j}\right)$
$=$ $\sum_{i}p_{i}\ell_{0}(q_{i})+(1-p_{i})\ell_{1}\left(1-q_{i}\right)$
$\geq$ $\sum_{i}p_{i}\ell_{0}(p_{i})+(1-p_{i})\ell_{1}\left(1-p_{i}\right)$
$=L(\bm{p},\bm{p}),$
[…]